OpenStudy (anonymous):

if f(x) = 2x + 3, f(c + 2) = 15, f(c) =

OpenStudy (anonymous):

\(f(x)=2x+3\) so \(f(c)=2c+3\) the variable is not important

OpenStudy (anonymous):

oh sorry i misread the question \(f(c+2)=2(c+2)+3=15\) we can solve for \(c\)

OpenStudy (anonymous):

\[2(c+2)+3=15\] \[2c+4+3=25\] \[2c+7=15\] \[2c=8\] \[c=4\] now that we know \(c=4\) we can find \(f(c)=f(4)\)

OpenStudy (anonymous):

would f(4) be 6?

hero (hero):

Hint: f(4) implies that x = 4

OpenStudy (anonymous):

@Hero so would it be 11?

hero (hero):

Yup

OpenStudy (anonymous):

thanks

hero (hero):

You're the first person who has successfully been able to make sense of my hint for functions of this type. Consider yourself competent :D

OpenStudy (anonymous):

:D